OpenStudy (anonymous):

Solve for x: 4x divided by x-3 = 2 + 12 divided by x-3 ...?

OpenStudy (watchmath):

Is that correct that the numerator of the right side is 2+12? or is that 2x+12 ?

OpenStudy (anonymous):

yes it's correct. but the 2 is not part of the numerator..just the 12

OpenStudy (watchmath):

So is it like this? \(\frac{4x}{x-3}=2+\frac{12}{x-3}\)

OpenStudy (anonymous):

yes.

OpenStudy (watchmath):

Ok, multiply the equation by \(x-3\) we get \(4x=2(x-3)+12\) \(4x=2x-6+12\) \(2x=6\) \(x=3\) But if we plug in back \(x=3\) to the original equation we are in trouble because we divide by zero. So the equation has no solution!.

OpenStudy (anonymous):

Thanks!